OpenStudy (anonymous):

a store is having a 10% off sale. it gives a additional 20% off if the item is still above 100 dollars after the first discount. Jacket costs 124.99, if you get the second discount how many dollars have you saved in total?

OpenStudy (anonymous):

\[124.99\times .9\times .8=89.99\] for a savings of 124.99 - 89.99 = $35